OpenStudy (konradzuse):

Find the equation for the funciton f that has the given derivative and whose graph passes through the given point.

OpenStudy (konradzuse):

f'(x) = 2 sin (4x) Point is \[(\pi/4, -17/2)\]

OpenStudy (anonymous):

Hm, integrate the function?

OpenStudy (anonymous):

good idea then find C by replacement

OpenStudy (anonymous):

Yes, since \(x\) and \(y\) are given.

OpenStudy (konradzuse):

well integrating would be -2cos(4x) + c right?

OpenStudy (anonymous):

f(x)=-0.5cos(4x)+C and sub that point into it

OpenStudy (anonymous):

then u can find C

OpenStudy (konradzuse):

ah okay

OpenStudy (konradzuse):

\[\frac{-\cos(4(\pi/4))}{2} \] which would be \[\frac{-\cos(4\pi/16)}{2}\]???

OpenStudy (konradzuse):

\[\frac{-\cos4(−17/2)}{2} = \frac{-\cos4(−68/8)}{2} \] ???

OpenStudy (konradzuse):

err forget that 4 before the (-68/8) :(

OpenStudy (anonymous):

I think that you are confused; one value is \(x\) and the other is \(y\) you can't sub it into the same thing.

OpenStudy (konradzuse):

Yeah I'm konfused, Idk what I'm doing :(

OpenStudy (konradzuse):

so I ijust use the first point then? pi/4

OpenStudy (anonymous):

\[-\frac{17}{2} =-0.5\cos\left(4*\frac{\pi}{4}\right) +C\]

OpenStudy (anonymous):

Solve for \(C\)

OpenStudy (anonymous):

\[2\int sin(4x)=\frac{1}{2}\cos(4x)+C\] \[f(x)=\frac{1}{2}\cos(4x)+C\] \[f(\frac{\pi}{4})=\frac{1}{2}\cos(2\pi)+C=-\frac{17}{2}\]

OpenStudy (anonymous):

\[\frac{1}{2}+C=-\frac{17}{2}\] \[C=\frac{-17}{2}-\frac{1}{2}=-9\] if i did not mess up

OpenStudy (anonymous):

-0.5cos(4*pi/4)+C=-17/2 that is -0.5cos(pi)+C=-17/2 -0.5cos(pi)=0.5 0.5+C=-17/2 C=-9

OpenStudy (anonymous):

ho ho ho i did mess up!

OpenStudy (anonymous):

twice

OpenStudy (konradzuse):

So the equation is 12cos(4x)+C or 12cos(4x)+ (-9)?

OpenStudy (konradzuse):

f(x) = that :P

OpenStudy (anonymous):

-0.5cos(4x)+(-9)=f(x)

OpenStudy (anonymous):

\[2\int\sin(4x)=-\frac{1}{2}\cos(4x)+C\] and \[f(\frac{\pi}{4})=-\frac{1}{2}\cos(\pi)+C=-\frac{17}{2}\] \[C=-9\]

OpenStudy (konradzuse):

idk why it said 12cos LOL :( stoopid copy/paste

OpenStudy (konradzuse):

Thanks all!

OpenStudy (anonymous):

i thk it is book's problem -.-